OpenStudy (mirandaregina):

How do I find the endpoint of a segment? I'm trying to follow an example that goes: A (-5,7) B (x2, y2) M(-2,2) Find the coordinates of B. -5+x2/2, 7+y2/2 = (-2,2) are the coordinates of B. My question is, how does -5+x2/2 = -2 and 7/y2 = 2??

OpenStudy (amistre64):

area A and M the endpoints?

OpenStudy (amistre64):

*are A and M ....

OpenStudy (mirandaregina):

@amistre64 No, A is the midpoint and M is one of the endpoints

OpenStudy (amistre64):

if A is the middle, the A is also the average of the end points ... does that make sense?

OpenStudy (amistre64):

do you know how to find an average?

OpenStudy (amistre64):

it appears from the information given , that A and B are endpoints, and M is the middle point between them to me

OpenStudy (mirandaregina):

@amistre64 you're right, A and B are the endpoints. Sorry about that! And no, I don't know how to find an average. I'm trying to help my little sister out with her math but I can't remember very much from Geometry here :/

OpenStudy (amistre64):

heres a long version that we can work into a formula: spose we start at point: A we need to find the distance (the difference) between A and B, the endpoints: B-A now half the distance is just (B-A)/2 and we can just add the to the start of it: A + (B-A)/2 now we can manipulate this into (2A + B - A) /2 (A + B) /2

OpenStudy (amistre64):

half way between 2 points, is simply the sum of the points, divided in half

OpenStudy (amistre64):

spose we have the points 3 and 15 (3+15)/2 = 18/2 = 9 9 is halfway between 2 and 15

OpenStudy (amistre64):

A (-5,7) B (x2, y2) M(-2,2) midpoint is halfway between the xs, ad halfway between the ys (ax + bx)/2 = mx (-5+bx)/2 = -2, solve for bx (ay + by)/2 = my (7 + by)/2 = 2, solve for by

OpenStudy (mirandaregina):

@amistre64 So, (-5 + bx)/2 = -2 is where I get confused because I get (-5 + bx)/2 = -5/2. Clearly that isn't right. How did you get -2?

OpenStudy (amistre64):

-2 is already stated as the x part of the point M: M = (-2,2) therefore the middle between A and B for the x parts has the value of -2

OpenStudy (amistre64):

(-5 + bx)/2 = -2 multiply both sides by 2 -5 + bx = -2(2) now add 5 to both sides bx = -2(2) + 5

OpenStudy (mirandaregina):

@amistre64 ok this is making more sense to me now lol I really appreciate you helping me to understand this!

OpenStudy (amistre64):

youre welcome